OpenStudy (anonymous):

The area of a square garden is 300m². How long is the diagonal? A. 5√6 m B. 150 m C. 10√6 m D.900m

OpenStudy (anonymous):

Area of a square: \[s ^{2}\] s = length of side ____________________________________ Formula for the diagonal: \[2s ^{2}\] ____________________________________ So, the diagonal would be: \[\sqrt{2 \times s ^{2}}\] \[\sqrt{600}\] \[10\sqrt{6 m}\]

OpenStudy (anonymous):

Oh, okay. Thanks a lot!